Find the equation (in terms of x) of the line through the points (-5,-4) and (2,1). y=?
Answer by rfadrogane(214)

Find the equation (in terms of x) of the line through the points (-5,-4) and (2,1).
Sol'n:
Solve for the slope first:
m =(y2-y2)/(x2-x1) let P1(-5,-4) & P2(2,1).
= (1+4)/(2+5)
m = 5/7
using point-slope form:
y-y1 = m(x-x1) @ P1(-5,-4)
y+4 = 5/7(x+5)
7(y+4)= 5(x+5)
7y+28 = 5x+25
5x-7y-3 = 0 ---answer